How many quarts of pure antifreeze must be added to 4 quarts of a 40% antifreeze solution to obtain a 60% antifreeze solution?
Answer by jim_thompson5910(35256)

"4 quarts of a 40% antifreeze solution" ---> 4*0.40 = 1.6 quarts of pure antifreeze.

Adding some unknown amount, call it x, of pure antifreeze gives you 1.6+x quarts of pure antifreeze.

-------------------------------------------------------

You have 4 quarts of a mixed solution and you're adding x quarts of pure antifreeze to the mix.
So you end up with a total of 4+x quarts of mixed solution.

-------------------------------------------------------

(amount of pure antifreeze)/(amount of solution) = percent of antifreeze


(1.6 + x)/(4 + x) = 0.60


1.6 + x = 0.60(4 + x)


1.6 + x = 0.60(4) + 0.60(x)


1.6 + x = 2.4 + 0.60x


1.6 + x - 0.60x = 2.4


x - 0.60x = 2.4 - 1.6


0.40x = 0.8


x = 0.8/0.40


x = 2


You need to add 2 quarts of pure antifreeze.